# The impact of $\gamma N$ and $\gamma^{*}N$ interactions on our understanding of nucleon excitations

### v1

@@ 6.3   Hybrid resonances and cascade decays
These possibilities can be explored by analyzing cascade decays, as studied by the CBELSA/TAPS collaboration [96, 1056]. Figure 3.9 shows the Dalitz plot for the reaction $\gamma p\to p\pi^{0}\pi^{0}$ in the invariant mass range from 1975 to 2225 MeV. Clear bands are observed, demonstrating the significance of cascade decays via intermediate isobars. The most prominent contributions arise from the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and $N(1680)\nicefrac{{5}}{{2}}^{+}$. Interestingly, Refs. [96, 1056] report markedly different decay patterns for the positive-parity $N^{*}$ and $\Delta^{*}$ resonances with masses around 1900 MeV. The four $N^{*}$ resonances $N(1880)\nicefrac{{1}}{{2}}^{+}$, $N(1965)\nicefrac{{3}}{{2}}^{+}$, $N(2000)\nicefrac{{5}}{{2}}^{+}$, $N(1990)\nicefrac{{7}}{{2}}^{+}$ likely form a 70-plet (cf. Fig. 3.2) and decay with an average branching fraction of $(34\pm 6)$into $N\pi$ and $\Delta\pi$, and with a branching fraction of $(21\pm 5)$into orbitally excited states such as $N(1520)\nicefrac{{3}}{{2}}^{-}\pi$, $N(1535)\nicefrac{{1}}{{2}}^{-}\pi$, and $N\sigma$. In contrast, the four $\Delta^{*}$ resonances $\Delta(1910)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1920)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$, $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ are members of a 56-plet and have an average decay branching fraction of $(44\pm 7)$into $N\pi$ or $\Delta\pi$. Their decay into orbitally excited states is significantly smaller with only $(5\pm 2)$[96]. This different behavior can be explained by their different quark-model wave functions: The four positive-parity $\Delta^{*}$ states have a symmetric spatial wave function $\phi_{n_{\rho}n_{\lambda},l_{\rho}l_{\lambda}}\sim\phi_{00,20}+\phi_{00,02}$, where either the $\rho$ or $\lambda$ oscillator is excited to $l_{\rho}=2$ or $l_{\lambda}=2$, so these states can release their excitation energy in a single step. The four positive-parity $N^{*}$ states in the 70-plet have a mixed-symmetric component $\phi_{00,20}-\phi_{00,02}$, but also a mixed-antisymmetric component $\phi_{00,11}$ where both oscillators are excited to $l_{\rho}=l_{\lambda}=1$. If one of the oscillators de-excites, the other remains excited; thus, a second de-excitation step is needed, as illustrated in Fig. 3.9, which may lead to the observed cascade decays.
@@
On the other hand, the Bonn model of baryon resonances predicts the lowest-mass $J^{P}=\nicefrac{{1}}{{2}}^{+}$ state in the fourth excitation shell to be “Roper-like” with a mass of 2009 MeV [549], making it a plausible candidate for an identification with the $N(2100)\nicefrac{{1}}{{2}}^{+}$. Its wave function should then contain a large contribution from the harmonic-oscillator wave function
@@
Finally, if the $N(2100)\nicefrac{{1}}{{2}}^{+}$ is a hybrid, it likely decays via the breaking of the excited string, leading to intrinsic orbital angular momentum in one of the decay products. This decay mechanism was originally proposed for mesons [1057, 1058] and is likely applicable to the decays of hybrid baryons as well. Indeed, for the $\pi_{1}(1600)$, the channels $f_{1}(1285)\pi$ [1059] and $b_{1}(1235)\pi$ [1060] are observed to be strong decay modes, although $\rho\pi$ [1061], $\eta^{\prime}\pi$, and $f_{1}(1285)\pi$ [1059] are established decay channels as well. We may thus expect some two-body decays, but a significant fraction should proceed into orbitally excited states; in particular the $S$-wave decay into $N(1535)\nicefrac{{1}}{{2}}^{-}$ and a pion should be strong. Radially excited states are not expected to appear in such cascade decays.
@@ 6.4   Diquark models and “missing resonances”
The similarity of the Regge trajectories suggests a quark-diquark structure of baryon resonances. There are different interpretations of the term diquark. In models with a frozen diquark, only excitations involving a quasi-stable diquark are allowed. In $SU(6)$, positive-parity diquark excitations belong to a 56-plet, while negative-parity diquark excitations belong to a 70-plet. Baryons with $P=+1$ in a 70-plet indicate broken symmetry [661]. States belonging to the 20-plet are strictly forbidden in frozen diquark models. The role of diquarks in spectroscopy, in baryon decays, and in the reaction dynamics of baryons is reviewed in Ref. [520] and, more recently, in Ref. [522].
@@ 6.5   Chiral symmetry restoration
Glozman observed that in the baryon spectrum, especially at higher masses, approximate parity doublets appear [1062, 1063]. He pointed out several pairs of nucleon and $\Delta$ resonances with the same spin but opposite parity and nearly degenerate masses, such as:
@@
Subsequent discoveries of new resonances have lent further support to this idea. For example, the $N(1900)\nicefrac{{3}}{{2}}^{+}$ is close in mass to the $N(1875)\nicefrac{{3}}{{2}}^{-}$, and the mass of the $N(1880)\nicefrac{{1}}{{2}}^{+}$ is close to that of the $N(1895)\nicefrac{{1}}{{2}}^{-}$. The concept of parity doubling was later extended to the meson sector as well [1064]. In a comprehensive review [1065], the idea of chiral symmetry restoration was further expanded to include partial restoration of $U(1)_{A}$ symmetry. In this scenario, not only are positive- and negative-parity states expected to become approximately degenerate at high masses, but also $N^{*}$ and $\Delta^{*}$ resonances with the same spin $J$ may converge, reflecting a higher degree of symmetry in the baryon spectrum.
Figure 6.7: (color online) Squared masses of $\Delta^{*}$ resonances with $J=L+S$ (stretched states) in comparison to squared meson masses [1066]. With kind permission of The European Physical Journal (EPJ).
Figure 6.8: The pseudo-$\chi^{2}$ of the fit to a large body of pion- and photo-produced reactions as a function of the mass of $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ (solid points) or $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ (open circles). The scale on the left (right) abscissa refers to the $\nicefrac{{7}}{{2}}^{+}$ ($\nicefrac{{7}}{{2}}^{-}$) partial wave. The curves are to guide the eye. The Figure is taken from [417].
@@
This interpretation of parity partners as evidence for chiral symmetry restoration was questioned in Ref. [1067]. An empirical mass formula for baryon resonances [1068] suggested that the masses of light baryon resonances do not follow a $M^{2}\propto l+2n$ behavior, but rather $M\propto l+n$ (where $l=l_{\rho}+l_{\lambda}$ and $n=n_{\rho}+n_{\lambda}$). Thus, the quartet of positive-parity states $\Delta(1910)\nicefrac{{1}}{{2}}^{+}\dots\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, with a leading $l=2$ and $n=0$ contribution, and the triplet of negative-parity states $\Delta(1900)\nicefrac{{1}}{{2}}^{-},\Delta(1930)\nicefrac{{3}}{{2}}^{-},\Delta(1940)\nicefrac{{5}}{{2}}^{-}$, with leading $l=1$ and $n=1$, are expected to be mass-degenerate. However, in this scheme, no $J^{P}=\nicefrac{{7}}{{2}}^{-}$ state is expected at this mass. Such a state would require $l=3$ and is expected at about 2150 MeV. At the time, the $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ had a 1* rating, and uncertainties of about 80 MeV were assigned to the mass determination. Also, further resonances with the predicted properties could be missing (like in the case of $N(1900)\nicefrac{{3}}{{2}}^{+}$ where Glozman suggested the $N(2080)\nicefrac{{3}}{{2}}^{-}$ as parity partner, while later the $N(1875)\nicefrac{{3}}{{2}}^{-}$ – in principle a perfect parity partner – was discovered [417]).
Data on photoproduction enabled a more refined study of the $I=3/2$, $J^{P}=\nicefrac{{7}}{{2}}^{-}$ channel. In a coupled-channel analysis [166], the masses of possible $\Delta^{*}$ resonances with $J^{P}=\nicefrac{{7}}{{2}}^{+}$ and $J^{P}=\nicefrac{{7}}{{2}}^{-}$ were scanned over a wide mass range. The results of a fit to an extensive data set on pion- and photo-induced reactions are shown in Fig. 6.8. The scans reveal clear minima, from which Breit-Wigner masses of $(1917\pm 4)$ and $(2176\pm 40)$ MeV were derived. The small coupling of the $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ to $\pi N$ (branching fraction $(3.3\pm 1.5)$%) explains why it was so difficult to observe it in $\pi N$ elastic scattering. The observed $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ mass supports quark models and AdS/QCD, but conflicts with models in which chiral symmetry is restored in meson and baryon resonances below 2 GeV. In summary, while these experimental findings are not in conflict with the concept of parity doublets in general, they disfavor their identification with the low-lying excitations in the baryon spectrum.
@@ 7   Summary and Outlook
It is worth noting that in the late 1970s, $\Delta^{*}$ resonances were reported with masses as high as 4.1 GeV, spins up to $J=\nicefrac{{21}}{{2}}$ and orbital angular momenta up to $L=9$ [1069]. In the RPP, these states are listed under the $\Delta(3000)$ region. Relative to the known Regge trajectory of $\Delta^{*}$s, these high-spin resonances have masses that are too high. Can the gluon string connecting a quark and a diquark persist over such large distances for long enough to sustain a resonance? If so, does the associated Regge trajectory begin to curve in the far-distance confinement region? Investigating this would likely require a $\pi^{+}$ beam — to produce $\Delta^{++}$ baryons via interactions with a proton target — with momenta reaching up to 10 GeV.
@@
Effective field theories (EFTs) have become a powerful tool for interpreting new data. Rather than resolving the quark substructure of baryons, they focus on the dynamics of hadronic interactions between mesons and baryons. In fitting the data, some resonance poles must be included explicitly, while others emerge dynamically from the interactions. Dynamically generated states are often considered unrelated to $qqq$ spectroscopy. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$, for instance, was long thought to be a dynamically generated state with no three-quark core. However, electroproduction experiments have resolved this issue, confirming that the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is indeed a genuine three-quark excited state at its core, albeit with a substantial meson cloud. The so-called Weinberg criterion — based on meson-baryon scattering amplitudes at zero momentum transfer — provides a way to estimate the degree to which hadrons are composite. Based on this criterium, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is characterized as a dynamically generated resonance with a notable three-quark core, and electroproduction data and EFTs now agree that the Roper resonance has a $qqq$ core and a significant meson cloud. In other cases, results remain contradictory. For example, the $N(1520)\nicefrac{{3}}{{2}}^{-}$ is sometimes described as a predominantly composite meson-baryon state, yet electroproduction data prove it has a clear three-quark structure with little meson cloud involvement.
@@
Several accelerators and colliders are still in operation, and new insight into the nucleon, its structure and its excitations are expected. At the low-energy end, electron-proton scattering experiments are in the planning at MESA in Mainz, Germany, aiming at precise measurements of electroweak coupling constants like the Weinberg angle, and searches for light dark matter will be pursued [1070]. The role of gluons in hadron spectroscopy has been a central topic since the advent of QCD. Gluonic degrees of freedom represent a challenging and often debated aspect of meson spectroscopy, while baryonic hybrids are even more complex. If hybrid baryons indeed have high masses, as suggested by lattice calculations, mixing with conventional $qqq$ states may occur. However, the presence of a heavy constituent gluon could suppress such mixing. At JLab, an energy upgrade to 22 GeV is planned [1071] to perform high-statistics electroproduction experiments over a broad range of momentum transfers $Q^{2}$ providing crucial insights into gluonic components within baryon resonance wave functions. The considerable higher $Q^{2}$ range that is achievable at this energy will significantly benefit the nucleon resonance program in probing the emerging hadron mass program [19]. In addition, cascade decays could help illuminate this issue. At high $Q^{2}$, backgrounds from $t$- and $u$-channel exchanges are reduced, offering a cleaner environment for future studies of the baryon resonance spectrum.
@@
At high energies, structure functions will remain a topical theme at JLab, at the electron-ion collider at the Brookhaven National Laboratory presently in the construction stage, and at the planned electron-ion collider in China. Gravity-like interactions are still in a very early stage but show great promise as a novel way of probing the strong interaction in hadrons and light nuclei. Gravity as a probe was theoretically studied already in the 1960’s [1072, 1073] but remained dormant for four decades. Maxim Polyakov [1074] considered using deeply virtual Compton scattering as a way to mimic gravity and probe the gravitational structure of the strong force. Following the first successful experimental application of this technique in 2018 with the ground-state proton, the field developed rapidly and has expanded to probe gravitational properties of nucleon resonances through the energy-momentum tensor [1075, 1076, 1077].

### v4

@@ 6.3  Hybrid resonances and cascade decays
These possibilities can be explored by analyzing cascade decays, as studied by the CBELSA/TAPS collaboration CBELSATAPS:2022uad , CBELSATAPS:2014wvh . Figure 3.9 shows the Dalitz plot for the reaction $\gamma p\to p\pi^{0}\pi^{0}$ in the invariant mass range from 1975 to 2225 MeV. Clear bands are observed, demonstrating the significance of cascade decays via intermediate isobars. The most prominent contributions arise from the $\Delta(1232)\nicefrac{{3}}{{2}}^{+}$, $N(1520)\nicefrac{{3}}{{2}}^{-}$, and $N(1680)\nicefrac{{5}}{{2}}^{+}$. Interestingly, Refs. CBELSATAPS:2022uad , CBELSATAPS:2014wvh report markedly different decay patterns for the positive-parity $N^{*}$ and $\Delta^{*}$ resonances with masses around 1900 MeV. The four $N^{*}$ resonances $N(1880)\nicefrac{{1}}{{2}}^{+}$, $N(1965)\nicefrac{{3}}{{2}}^{+}$, $N(2000)\nicefrac{{5}}{{2}}^{+}$, $N(1990)\nicefrac{{7}}{{2}}^{+}$ likely form a 70-plet (cf. Fig. 3.2) and decay with an average branching fraction of $(34\pm 6)$% into $N\pi$ and $\Delta\pi$, and with a branching fraction of $(21\pm 5)$% into orbitally excited states such as $N(1520)\nicefrac{{3}}{{2}}^{-}\pi$, $N(1535)\nicefrac{{1}}{{2}}^{-}\pi$, and $N\sigma$. In contrast, the four $\Delta^{*}$ resonances $\Delta(1910)\nicefrac{{1}}{{2}}^{+}$, $\Delta(1920)\nicefrac{{3}}{{2}}^{+}$, $\Delta(1905)\nicefrac{{5}}{{2}}^{+}$, $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ are members of a 56-plet and have an average decay branching fraction of $(44\pm 7)$% into $N\pi$ or $\Delta\pi$. Their decay into orbitally excited states is significantly smaller with only $(5\pm 2)$CBELSATAPS:2022uad . This different behavior can be explained by their different quark-model wave functions: The four positive-parity $\Delta^{*}$ states have a symmetric spatial wave function $\phi_{n_{\rho}n_{\lambda},l_{\rho}l_{\lambda}}\sim\phi_{00,20}+\phi_{00,02}$, where either the $\rho$ or $\lambda$ oscillator is excited to $l_{\rho}=2$ or $l_{\lambda}=2$, so these states can release their excitation energy in a single step. The four positive-parity $N^{*}$ states in the 70-plet have a mixed-symmetric component $\phi_{00,20}-\phi_{00,02}$, but also a mixed-antisymmetric component $\phi_{00,11}$ where both oscillators are excited to $l_{\rho}=l_{\lambda}=1$. If one of the oscillators de-excites, the other remains excited; thus, a second de-excitation step is needed, as illustrated in Fig. 3.9, which may lead to the observed cascade decays.
@@
On the other hand, the Bonn model of baryon resonances predicts the lowest-mass $J^{P}=\nicefrac{{1}}{{2}}^{+}$ state in the fourth excitation shell to be “Roper-like” with a mass of 2009 MeV Loring:2001kx , making it a plausible candidate for an identification with the $N(2100)\nicefrac{{1}}{{2}}^{+}$. Its wave function should then contain a large contribution from the harmonic-oscillator wave function
@@
Finally, if the $N(2100)\nicefrac{{1}}{{2}}^{+}$ is a hybrid, it likely decays via the breaking of the excited string, leading to intrinsic orbital angular momentum in one of the decay products. This decay mechanism was originally proposed for mesons Isgur:1985vy , Kokoski:1985is and is likely applicable to the decays of hybrid baryons as well. Indeed, for the $\pi_{1}(1600)$, the channels $f_{1}(1285)\pi$ E852:2004gpn and $b_{1}(1235)\pi$ E852:2004rfa are observed to be strong decay modes, although $\rho\pi$ COMPASS:2018uzl , $\eta^{\prime}\pi$, and $f_{1}(1285)\pi$ E852:2004gpn are established decay channels as well. We may thus expect some two-body decays, but a significant fraction should proceed into orbitally excited states; in particular the $S$-wave decay into $N(1535)\nicefrac{{1}}{{2}}^{-}$ and a pion should be strong. Radially excited states are not expected to appear in such cascade decays.
@@ 6.4  Diquark models and “missing resonances”
The similarity of the Regge trajectories suggests a quark-diquark structure of baryon resonances. There are different interpretations of the term diquark. In models with a frozen diquark, only excitations involving a quasi-stable diquark are allowed. In $SU(6)$, positive-parity diquark excitations belong to a 56-plet, while negative-parity diquark excitations belong to a 70-plet. Baryons with $P=+1$ in a 70-plet indicate broken symmetry Lichtenberg:1969sxc . States belonging to the 20-plet are strictly forbidden in frozen diquark models. The role of diquarks in spectroscopy, in baryon decays, and in the reaction dynamics of baryons is reviewed in Ref. Anselmino:1992vg and, more recently, in Ref. Barabanov:2020jvn .
@@ 6.5  Chiral symmetry restoration
Glozman observed that in the baryon spectrum, especially at higher masses, approximate parity doublets appear Glozman:1999tk , Glozman:2002kq . He pointed out several pairs of nucleon and $\Delta$ resonances with the same spin but opposite parity and nearly degenerate masses, such as:
@@
Subsequent discoveries of new resonances have lent further support to this idea. For example, the $N(1900)\nicefrac{{3}}{{2}}^{+}$ is close in mass to the $N(1875)\nicefrac{{3}}{{2}}^{-}$, and the mass of the $N(1880)\nicefrac{{1}}{{2}}^{+}$ is close to that of the $N(1895)\nicefrac{{1}}{{2}}^{-}$. The concept of parity doubling was later extended to the meson sector as well Glozman:2002jf . In a comprehensive review Glozman:2007ek , the idea of chiral symmetry restoration was further expanded to include partial restoration of $U(1)_{A}$ symmetry. In this scenario, not only are positive- and negative-parity states expected to become approximately degenerate at high masses, but also $N^{*}$ and $\Delta^{*}$ resonances with the same spin $J$ may converge, reflecting a higher degree of symmetry in the baryon spectrum.
Figure 6.7: (color online) Squared masses of $\Delta^{*}$ resonances with $J=L+S$ (stretched states) in comparison to squared meson masses Klempt:2012fy . With kind permission of The European Physical Journal (EPJ).
Figure 6.8: The pseudo-$\chi^{2}$ of the fit to a large body of pion- and photo-produced reactions as a function of the mass of $\Delta(1950)\nicefrac{{7}}{{2}}^{+}$ (solid points) or $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ (open circles). The scale on the left (right) abscissa refers to the $\nicefrac{{7}}{{2}}^{+}$ ($\nicefrac{{7}}{{2}}^{-}$) partial wave. The curves are to guide the eye. The Figure is taken from Anisovich:2011sv .
@@
This interpretation of parity partners as evidence for chiral symmetry restoration was questioned in Ref. Klempt:2002tt . An empirical mass formula for baryon resonances Klempt:2002vp suggested that the masses of light baryon resonances do not follow a $M^{2}\propto l+2n$ behavior, but rather $M\propto l+n$ (where $l=l_{\rho}+l_{\lambda}$ and $n=n_{\rho}+n_{\lambda}$). Thus, the quartet of positive-parity states $\Delta(1910)\nicefrac{{1}}{{2}}^{+}\dots\Delta(1950)\nicefrac{{7}}{{2}}^{+}$, with a leading $l=2$ and $n=0$ contribution, and the triplet of negative-parity states $\Delta(1900)\nicefrac{{1}}{{2}}^{-},\Delta(1930)\nicefrac{{3}}{{2}}^{-},\Delta(1940)\nicefrac{{5}}{{2}}^{-}$, with leading $l=1$ and $n=1$, are expected to be mass-degenerate. However, in this scheme, no $J^{P}=\nicefrac{{7}}{{2}}^{-}$ state is expected at this mass. Such a state would require $l=3$ and is expected at about 2150 MeV. At the time, the $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ had a 1* rating, and uncertainties of about 80 MeV were assigned to the mass determination. Also, further resonances with the predicted properties could be missing (like in the case of $N(1900)\nicefrac{{3}}{{2}}^{+}$ where Glozman suggested the $N(2080)\nicefrac{{3}}{{2}}^{-}$ as parity partner, while later the $N(1875)\nicefrac{{3}}{{2}}^{-}$ – in principle a perfect parity partner – was discovered Anisovich:2011sv ).
Data on photoproduction enabled a more refined study of the $I=3/2$, $J^{P}=\nicefrac{{7}}{{2}}^{-}$ channel. In a coupled-channel analysis Anisovich:2015gia , the masses of possible $\Delta^{*}$ resonances with $J^{P}=\nicefrac{{7}}{{2}}^{+}$ and $J^{P}=\nicefrac{{7}}{{2}}^{-}$ were scanned over a wide mass range. The results of a fit to an extensive data set on pion- and photo-induced reactions are shown in Fig. 6.8. The scans reveal clear minima, from which Breit-Wigner masses of $(1917\pm 4)$ and $(2176\pm 40)$ MeV were derived. The small coupling of the $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ to $\pi N$ (branching fraction $(3.3\pm 1.5)$%) explains why it was so difficult to observe it in $\pi N$ elastic scattering. The observed $\Delta(2200)\nicefrac{{7}}{{2}}^{-}$ mass supports quark models and AdS/QCD, but conflicts with models in which chiral symmetry is restored in meson and baryon resonances below 2 GeV. In summary, while these experimental findings are not in conflict with the concept of parity doublets in general, they disfavor their identification with the low-lying excitations in the baryon spectrum.
@@ 7  Summary and Outlook
It is worth noting that in the late 1970s, $\Delta^{*}$ resonances were reported with masses as high as 4.1 GeV, spins up to $J=\nicefrac{{21}}{{2}}$ and orbital angular momenta up to $L=9$ Hendry:1978cd . In the RPP, these states are listed under the $\Delta(3000)$ region. Relative to the known Regge trajectory of $\Delta^{*}$s, these high-spin resonances have masses that are too high. Can the gluon string connecting a quark and a diquark persist over such large distances for long enough to sustain a resonance? If so, does the associated Regge trajectory begin to curve in the far-distance confinement region? Investigating this would likely require a $\pi^{+}$ beam — to produce $\Delta^{++}$ baryons via interactions with a proton target — with momenta reaching up to 10 GeV.
@@
Effective field theories (EFTs) have become a powerful tool for interpreting new data. Rather than resolving the quark substructure of baryons, they focus on the dynamics of hadronic interactions between mesons and baryons. In fitting the data, some resonance poles must be included explicitly, while others emerge dynamically from the interactions. Dynamically generated states are often considered unrelated to qqqqqq spectroscopy. The Roper resonance $N(1440)\nicefrac{{1}}{{2}}^{+}$, for instance, was long thought to be a dynamically generated state with no three-quark core. However, electroproduction experiments have resolved this issue, confirming that the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is indeed a genuine three-quark excited state at its core, albeit with a substantial meson cloud. The so-called Weinberg criterion — based on meson-baryon scattering amplitudes at zero momentum transfer — provides a way to estimate the degree to which hadrons are composite. Based on this criterium, the $N(1440)\nicefrac{{1}}{{2}}^{+}$ is characterized as a dynamically generated resonance with a notable three-quark core, and electroproduction data and EFTs now agree that the Roper resonance has a qqqqqq core and a significant meson cloud. In other cases, results remain contradictory. For example, the $N(1520)\nicefrac{{3}}{{2}}^{-}$ is sometimes described as a predominantly composite meson-baryon state, yet electroproduction data prove it has a clear three-quark structure with little meson cloud involvement.
@@
Several accelerators and colliders are still in operation, and new insight into the nucleon, its structure and its excitations are expected. The doubling of the CEBAF accelerator from 6 GeV to 12 GeV maximum energy has opened up the kinematic space in hadron physics significantly. It extends both the mass reach in the photoproduction of excited baryons, and the reach in photon virtuality $Q^{2}$ in the electroproduction of mesons. The former program may be part of the GlueX experiment in Hall D  GlueX:2020idb , while the latter enables probing the excited baryons at significantly shorter distances and should lead to more quantifiable results in identifying the effective degrees of freedom. This is a focus of the newly constructed CLAS12 spectrometer Burkert:2020akg in Hall B. Much of the data that may be of interest to readers of this review have already been taken and are in the analysis stage. Another program that can be more rigorously pursued with CLAS12 is the study of nucleon mass creation using excited baryon states Achenbach:2025kfx .
At the low-energy end, electron-proton scattering experiments are in the planning at MESA in Mainz, Germany, aiming at precise measurements of electroweak coupling constants like the Weinberg angle, and searches for light dark matter will be pursued Schlimme:2024eky . The role of gluons in hadron spectroscopy has been a central topic since the advent of QCD. Gluonic degrees of freedom represent a challenging and often debated aspect of meson spectroscopy, while baryonic hybrids are even more complex. If hybrid baryons indeed have high masses, as suggested by lattice calculations, mixing with conventional qqqqqq states may occur. However, the presence of a heavy constituent gluon could suppress such mixing. At JLab, an energy upgrade to 22 GeV is planned Accardi:2023chb to perform high-statistics electroproduction experiments over a broad range of momentum transfers $Q^{2}$ providing crucial insights into gluonic components within baryon resonance wave functions. The considerable higher $Q^{2}$ range that is achievable at this energy will significantly benefit the nucleon resonance program in probing the emerging hadron mass program Achenbach:2025kfx . In addition, cascade decays could help illuminate this issue. At high $Q^{2}$, backgrounds from $t$- and $u$-channel exchanges are reduced, offering a cleaner environment for future studies of the baryon resonance spectrum.
@@
At high energies, structure functions will remain a topical theme at JLab, at the electron-ion collider at the Brookhaven National Laboratory presently in the construction stage, and at the planned electron-ion collider in China. Gravity-like interactions are still in a very early stage but show great promise as a novel way of probing the strong interaction in hadrons and light nuclei. Gravity as a probe was theoretically studied already in the 1960’s Kobzarev:1962wt , Pagels:1966zza but remained dormant for four decades. Maxim Polyakov Polyakov:2002yz considered using deeply virtual Compton scattering as a way to mimic gravity and probe the gravitational structure of the strong force. Following the first successful experimental application of this technique in 2018 with the ground-state proton, the field developed rapidly and has expanded to probe gravitational properties of nucleon resonances through the energy-momentum tensor Kim:2024hhd , Azizi:2020jog , Goharipour:2024atx .